### v1

@@ 5 Conclusions
In this study, we investigated a high-throughput 3D shape completion network for its applicability for estimating the volume of potato tubers on an operational harvester. Our research revealed that latent size 32 had the best 3D shape completion result. With that latent size, our CoRe++ network had an RMSE of 22.6 ml on the volumetric estimate, and this was better than the RMSE of the linear regression (31.1 ml) and the original CoRe network (36.9 ml). We also found that the RMSE of CoRe++ could be further reduced to 18.2 ml when performing the 3D shape completion in the center of the RGB-D image. With an average analysis time of 10 milliseconds per potato tuber, CoRe++ enables a high-throughput 3D shape completion up to 100 potato tubers per second. Hence, we can conclude that our network is able to quickly and accurately estimate the volume of fast-moving potato tubers on an operational harvester. Future improvements lie in expanding the dataset, testing on more potato cultivars, implementing a transformer-based network, and implementing rotating rollers on the harvester’s conveyor belt. We encourage other researchers to reuse our dataset and software.

### v3

@@ 5 Conclusions
In this study, we investigated a high-throughput 3D shape completion network for its suitability for estimating the volume of potato tubers on an operational harvester. Our research revealed that latent size 32 had the best 3D shape completion result. With that latent size, our CoRe++ network had an RMSE of 22.6 ml on the volumetric estimate, and this was better than the RMSE of the linear regression (31.1 ml) and the original CoRe network (36.9 ml). We also found that the RMSE of CoRe++ could be further reduced to 18.2 ml when performing the 3D shape completion in the center of the RGB-D image. With an average analysis time of 10 milliseconds per potato tuber, CoRe++ enables a high-throughput 3D shape completion up to 100 potato tubers per second. Hence, we can conclude that our network is able to quickly and accurately estimate the volume of fast-moving potato tubers on an operational harvester. Our method provides real-time potato tuber yield estimates, giving farmers valuable insights to further optimize planting, fertilization, and harvest strategies in subsequent growing seasons. This can help to increase yield, reduce waste, and better meet market demands—effects not easily achieved by current analytics. The continual progress in 3D shape completion and machine learning algorithms will enable more accurate and more useful yield monitoring, as well as mitigating the effects of occlusion and partial visibility in robotic harvesting, quality control and sorting.